OpenStudy (campbell_st):

y = (10 + 5x)/2 or y = 5/2 x + 5

OpenStudy (anonymous):

Goal is to get Y by itself add term 5x to both sides divide all terms by -2 -5x - 2y = -10 +5x +5x -2y = 5x -10 divide all terms by -2 y = -5/2 x +5

Directrix (directrix):

@ Carley --> See above for correct answer: y = -5/2 x +5